joe has twice as many 10 dollar bills as 5 dollar bills. he has a total of $500. how many % dollar bills does joe have? 5 dollar bills.
Answer by mananth(16946)

joe has twice as many 10 dollar bills as 5 dollar bills. he has a total of $500. how many % dollar bills does joe have? 5 dollar bills.
...
let $ 5 bills be x
$10 bills = 2x
..
5x+10*2x=500
5x+20x=500
25x=500
/25
x= 500/25
x=20
$5 20 bills
$10 40 bills
...
m.ananth@hotmail.ca